OpenStudy (goformit100):

Find the rate of change of the area of a circle with respect to its radius r when (a) r = 3 cm (b) r = 4 cm

OpenStudy (goformit100):

@terenzreignz

terenzreignz (terenzreignz):

\[\huge A = \pi r^2\]\[\huge \frac{dA}{dr}=2\pi r\] and just plug in.

OpenStudy (goformit100):

Thankyou Sir